OpenStudy (samigupta8):

If the binding energy per nucleon in Li and He nuclei are 5.60 MeV and 7.06 MeV respectively then in the reaction: p+ Li--------->2 He Energy of proton must be

OpenStudy (samigupta8):

@michele_laino sir plss..help

OpenStudy (michele_laino):

do we know the energy or the so called \(Q\) of such reaction?

OpenStudy (samigupta8):

Nope it is not mentioned....

OpenStudy (michele_laino):

please wait I make some computations...

OpenStudy (samigupta8):

Okk...i will...

OpenStudy (michele_laino):

if I use the connservation of the four-momentum, I get this: \[\Large {E_p} = \frac{{4{{\left( {{M_{He}}{c^2}} \right)}^2} - {{\left( {{M_p}{c^2}} \right)}^2} - {{\left( {{M_{Li}}{c^2}} \right)}^2}}}{{2{M_{Li}}{c^2}}}\] where \(E_p\) is the total energy of the proton

OpenStudy (samigupta8):

Sorry sir...bt i didn't get u...this formula From where it comes..??

OpenStudy (michele_laino):

it comes from the conservation of the four-momentum, namely we have to consider the vectors like objects of the Minkowski space

OpenStudy (samigupta8):

Four-momentum???

OpenStudy (michele_laino):

please what are your options?

OpenStudy (samigupta8):

28.24 17.28 1.46 39.2 all r in MeV

OpenStudy (michele_laino):

I'm thinking...

OpenStudy (samigupta8):

Okk...

OpenStudy (michele_laino):

my formula above is wrong

OpenStudy (samigupta8):

So what now ?

OpenStudy (michele_laino):

please wait...

OpenStudy (samigupta8):

So sorry....sir.

OpenStudy (michele_laino):

no problem :) the product of such reaction are two atoms of Helium?

OpenStudy (samigupta8):

Yep...

OpenStudy (michele_laino):

the number of mass is not conserved: \[\Large _1^1{\text{p}} + _3^6{\text{Li}} \to _2^4{\text{He}} + _2^4{\text{He}}\]

OpenStudy (samigupta8):

Lithium's isotope of mass number 7

OpenStudy (michele_laino):

as we can see at the left we have \(1+6=7\), whereas at the right we hae \(4+4=8\) ok!

OpenStudy (samigupta8):

Sorry didn't mention in the ques itself. ...

OpenStudy (michele_laino):

then we have a reaction like this: \[\Large _1^1{\text{p}} + _3^7{\text{Li}} \to _2^4{\text{He}} + _2^4{\text{He}}\]

OpenStudy (samigupta8):

Yo exactly

OpenStudy (michele_laino):

I think that we can apply the conservation of energy like in the photoelectric effect, so we can write: \[\huge E + 5.60 = 7.06\]

OpenStudy (samigupta8):

Bt there is some release in energy whenever a nuxlear reaction takes place

OpenStudy (samigupta8):

How can we apply energy conservation here?

OpenStudy (michele_laino):

yes! sure there is the so called \(Q\) of the reaction

OpenStudy (michele_laino):

I think that \(E\) stands for kinetic energy only of the proton

OpenStudy (samigupta8):

Then here no Q is mentioned then how can v proceed to the problem

OpenStudy (samigupta8):

Sorry sir it's 2:17 am here.....i need to go... Feeling very sleepy...plss..wud u help me vid dis problem l8r....

OpenStudy (michele_laino):

the only idea that I have in my mind is the equation above. Otherwise we have to consider the conservation of four-momentum, nevertheless, as I wrote before such conservation gives a wrong equation, since don't know the \(Q\) of the reaction

OpenStudy (michele_laino):

ok! I will help you!

OpenStudy (michele_laino):

given the reaction above, we can write this energy balance: \[\Large {M_p}{c^2} + {M_{Li}}{c^2} = 2{M_{He}}{c^2} + Q\] where \(Q\) is the energy freed by such reaction like gamma rays, otherwise such reaction is not possible

OpenStudy (samigupta8):

Yess correct we can write that since Q value is equal to the (mass of reactants-mass of product)*C^2

OpenStudy (samigupta8):

Bt sir how can we assume Q value to be 0 unless mentioned....?

OpenStudy (michele_laino):

after a substitution, we get: \[\Large 940 + 5.6 = \left( {2 \times 7.06} \right) + Q\] and we are able to compute the value of \(Q\)

OpenStudy (samigupta8):

What is 940 ??

OpenStudy (michele_laino):

the mass of proton, is: \[\Large {M_p}{c^2} = 940\;{\text{MeV}}\]

OpenStudy (samigupta8):

Okk 931 MeV r u saying that???

OpenStudy (samigupta8):

1 amu=931 MeV

OpenStudy (michele_laino):

yes! I have rounded the value

OpenStudy (samigupta8):

Ohh well...it's alryt...then...so do u want to say that energy of proton is 931 MeV in this question?

OpenStudy (michele_laino):

yes!

OpenStudy (samigupta8):

I mean is the ans according to u 940 MeV ?

OpenStudy (samigupta8):

Sorry bt that is not in the options ....

OpenStudy (michele_laino):

no it is not, since it is the energy balance in order to make such reaction

OpenStudy (samigupta8):

So after finding out the Q value of the reaction what shall we do ....n even v have calculated the q value by supposing the proton enrgy to be 940 MeV now how can we change it jst for the sake of ans...

OpenStudy (michele_laino):

please wait, I try to make some computations...

OpenStudy (samigupta8):

Okay sir!!!

OpenStudy (michele_laino):

I'm sorry I got a wrong result.

OpenStudy (michele_laino):

I go offline to review my notes on nuclear physics, and I try to solve such problem

OpenStudy (samigupta8):

It"zz okkay with me sir....

OpenStudy (anonymous):

1H + 7Li ® 4He + 4He B.E.(1)= 0.0 ± 0.0 keV B.E.(2)= 39244.526 ± 0.473 keV B.E.(3)= 28295.673 ± 0.005 keV B.E.(4)= 28295.673 ± 0.005 keV Q-value: 17346.82 keV Uncertainty: 0.473 keV (ignoring correlations) Threshold: 0 keV in order to make the reaction happen, we need energy of Q-value. I do not know whether 17.34 MeV correct. What is the answer?

OpenStudy (michele_laino):

here are my computations: the masses of nuclide are: \[\begin{gathered} {M_{Li}} = \left( {3 \times 938 + 4 \times 939} \right) - \left( {7 \times 5.6} \right) = 6530.8\;{\text{MeV}} \hfill \\ \hfill \\ {M_{He}} = \left( {2 \times 938 + 2 \times 939} \right) - \left( {4 \times 7.06} \right) = 3725.76\;{\text{MeV}} \hfill \\ \hfill \\ {M_p} = 938\;{\text{MeV}} \hfill \\ \hfill \\ {M_n} = 939\;{\text{MeV}} \hfill \\ \end{gathered} \] from the conservation of the four-momentum in the center of mass of products, the minimum kinetic energy of the proton, is: \[{E_p} = \frac{{4{{\left( {{M_{He}}} \right)}^2} - {{\left( {{M_p}} \right)}^2} - {{\left( {{M_{Li}}} \right)}^2}}}{{2{M_{Li}}}} \cong 917\;{\text{MeV}}\] and the \(Q\) of such reaction, is: \[938 + 6530.8 = 2 \times 3725.76 + Q \Rightarrow Q = 17.28\;{\text{MeV}}\]

OpenStudy (samigupta8):

@Michele_Laino sir the ans u r cuming up with as Q value here is the ans for the energy of proton....in the ans key...

OpenStudy (samigupta8):

Ans is 17.28 MeV

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (samigupta8):

Okk...sir bt we can use ur previous method also...where we can use direct formula for Q value as Binding energy of helium- binding energy of lithium ....

OpenStudy (michele_laino):

I think that both formulas are equivalent

OpenStudy (samigupta8):

From this we will get the Q value of the reaction n as u proposed to me that this Q value is what will be the kinetic energy of proton...so final ans is this only...

OpenStudy (michele_laino):

yes! Since I think that such \(Q\) value can be attributable like kinetic energy of the proton

OpenStudy (samigupta8):

Sir is it always the case ...like whatsoever is the enrgy released in a reaction would be attributed to some particles...

OpenStudy (michele_laino):

for example, let's suppose to consider this reaction: \[\Large \gamma \to {e^ + } + {e^ - }\] namely the pair creation

OpenStudy (michele_laino):

the minimum energy of the products is \(0.51 \times 2=1.02\) MeV, when electron and positron are at the rest

OpenStudy (michele_laino):

in order to make such reaction, the minimum energy of gamma ray, has to be just 1.02 MeV, otherwise such reaction doesn't happen

OpenStudy (samigupta8):

Okk so gamma particle will have to move with the energy mentioned in order to let go the reaction to its good pace....

OpenStudy (michele_laino):

yes!

OpenStudy (samigupta8):

Similarly in this case for reaction to occure we will ensure that the Q value goes to the enrgy of proton in order for completion of reaction....

OpenStudy (michele_laino):

here the rest mass of the electron or the rest mass of positron are 0.51 MeV

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (samigupta8):

Sir the Q value here will be negative in case of tge example u took....

OpenStudy (samigupta8):

The*

OpenStudy (michele_laino):

subsequently part of the total amount of energy goes for creation of the products namely the two atoms of helium, and the remaining part goes for gamma rays and recoil energy of the two atoms of helium

OpenStudy (samigupta8):

Bcoz the Q value = (mass of reactants- mass of products)*c^2 so will give us a negative value .. Isn't it...?

OpenStudy (michele_laino):

yes! this is not a contradiction. It means that part of the energy of the proton will be used in order to create the products

OpenStudy (samigupta8):

Okk so u mean that Q value will be 0 if we just take the energy of proton to be 17.28 MeV n that's how it has to be done...

OpenStudy (samigupta8):

Or better if we consider ur example only then energy of gamma particle should be such that Q value turns out to be just 0 Bcoz Q value -ve implies reaction will not occur

OpenStudy (michele_laino):

no, please I say that the \(Q\) value, is the minimum energy, kinetic energy since we have a free proton, which such proton have to have in order to make the reaction

OpenStudy (michele_laino):

I say free proton, since that proton is moving inside a space without any field of force

OpenStudy (samigupta8):

Sir was my statement wrong??? Why can't it be in that manner since we know that Q value has to be a positive quantity always for reaction to b feasible then we may consider the extreme case if it...n then solve the problem

OpenStudy (michele_laino):

total energy \(E\) of the proton, measured with respect to the laboratory system, is: \[\Large E = \left( {T + m} \right){c^2} = 17.28 + 938 = 955.28\;{\text{MeV}}\]

OpenStudy (michele_laino):

the \(Q\) of a reaction can be a negative number

OpenStudy (samigupta8):

BT then the reaction won't proceed

OpenStudy (michele_laino):

no, no, if we have a reaction with negative \Q\), it means that we have to add energy to the system

OpenStudy (michele_laino):

oops.. negative \(Q\)*...

OpenStudy (samigupta8):

Sorry sir ....It was given in one of the books I read....maybe it says that the reaction wouldn't be feasible thereby

OpenStudy (michele_laino):

I think that it is not spontaneous

OpenStudy (samigupta8):

Yupp.....that's what I meant

OpenStudy (samigupta8):

So u were saying something about conservation

OpenStudy (michele_laino):

yes! nevertheless if we add energy to the system, such reaction can be feasible

OpenStudy (samigupta8):

Feasible was the blunder .......lol

OpenStudy (michele_laino):

ok! :)

OpenStudy (michele_laino):

here is another example from my nuclear physics textbook

OpenStudy (samigupta8):

Yes ...I m here

OpenStudy (samigupta8):

U can carry on..

OpenStudy (michele_laino):

\[\huge n + p = d + 2.225\] neutron and proton are at the rest, and they combine in order to make a deuteron, with \(Q=2.225\) MeV